# Optimal Bias-Correction and Valid Inference in High-Dimensional Ridge Regression: A Closed-Form Solution

### v1

@@ 4   Empirical Application
where ${\mathbf{X}}_{t}$ consists of 122 macroeconomic variables, $y_{t}$ is the remaining target one, and ${\mathbf{f}}_{t}=(f_{1,t},...,f_{r,t})^{\prime}$ is an $r$-dimensional latent factor process which can be estimated by applying Principal Component Analysis (PCA) to ${\mathbf{X}}_{t}$. We set $q=10$ and $r=60$, and apply the ridge regression to estimate the parameters $\alpha_{i}$ and $\gamma_{j}$ for $1\leq i\leq q$ and $1\leq j\leq r$ with estimated factors. Hence, the number of covariates in the ridge regression is $p=q+r=70$. We focus on the prediction of the consumer price index: all (CPI-All), which is an important index related to the inflation. See also Stock and Watson (2002a) and Gao and Tsay (2022) for similar studies.
@@
First, we set $\lambda\in\{0.05n,0.1n,0.2n,...,1.5n\}$ which contains 16 candidates for the penalty parameter. We split the sample into two subsamples, where the first one consists of the first $80\%$ of the data for modeling and the second subsample of the remaining $20\%$ of the data for out-of-sample prediction. We adopt the rolling-window scheme as that in Gao and Tsay (2023), that is, we train the factors and the forecasting coefficients using the first subsample to predict the next target data point. Then we repeat the above procedure after moving the next available observation of predictors and target CPI-All variable from the second subsample to the first one to obtain the next prediction. This rolling-window scheme is terminated when there is no more observation to compute the forecasting error. We choose the optimal penalty in the ridge estimator such that the out-of-sample mean-squared forecast errors (MSFEs) achieve a minimum value. Thus, this setting is in line with the advocated models combining regularized estimation and data-driven selection of regularization parameters in Abadie and Kasy (2019). In this experiment, we find that the optimal penalty $\widehat{\lambda}=0.8n$ for $h=1$ and $\widehat{\lambda}=0.7n$ for $h=2$.
@@
Next, for each configuration of $(\widehat{\lambda},h)=(0.8n,1)$ and $(0.7n,2)$, we execute Algorithm 1. We observe that the bias-correction procedure terminates within 3 to 5 iterations when setting $\eta=O(10^{-4})$ in equation (24). For simplicity, we opt for $k=10$ in the bias-correction, a choice deemed sufficiently large. Following the guidelines provided in Section 2.7, we compute pointwise prediction intervals for the data points in the testing set. The standard errors for each prediction are determined using the method described in Remark 5(iv). We present the $95\%$ pointwise prediction intervals for both 1-step ahead and 2-step ahead forecasts in Figure 8 and Figure 9, respectively. It is evident that the majority of the true observations fall within these prediction intervals. Additionally, we evaluate the coverage rates of the prediction intervals across the 138 points in the testing set. We find a consistent coverage rate of $92.03\%$ for both the 1-step and 2-step ahead forecasts, indicating that the $95\%$ prediction intervals perform effectively in this empirical example.
@@
Figure 10: Out-of-sample pointwise prediction intervals for the monthly CPI-All from July 1, 2008 to Dec. 12, 2019, constructed using (36)-(37) at $95\%$ level with $\lambda^{*}=\lambda=0.5n$, $k=10$, $n^{*}=31$, and $l=10$ for the $1$-step ahead forecast. The blue lines represent the confidence intervals, while the dark lines plot the true observations.

### v2

@@ 4   An Empirical Application
where ${\mathbf{X}}_{t}$ consists of 122 macroeconomic variables, $y_{t}$ is the remaining target one, and ${\mathbf{f}}_{t}=(f_{1,t},...,f_{r,t})^{\prime}$ is an $r$-dimensional latent factor process which can be estimated by applying Principal Component Analysis (PCA) to ${\mathbf{X}}_{t}$. We set $q=10$ and $r=60$, and apply the ridge regression to estimate the parameters $\alpha_{i}$ and $\gamma_{j}$, for $1\leq i\leq q$ and $1\leq j\leq r$, with estimated factors. Hence, the number of covariates in the ridge regression is $p=q+r=70$. We focus on the prediction of the consumer price index: all (CPI-All), which is an important index related to the inflation. See also Stock and Watson (2002a) and Gao and Tsay (2022) for similar studies.
@@
First, we set $\lambda\in\{0.05n,0.1n,0.2n,...,1.5n\}$, which contains 16 candidates for the penalty parameter. We split the sample into two subsamples, where the first one consists of the first $80\%$ of the data for modelling and the second subsample of the remaining $20\%$ of the data for out-of-sample prediction. We adopt the rolling-window scheme as that in Gao and Tsay (2024), that is, we train the factors and the forecasting coefficients using the first subsample to predict the next target data point. Then we repeat the above procedure after moving the next available observation of predictors and target CPI-All variable from the second subsample to the first one to obtain the next prediction. This rolling-window scheme is terminated when there is no more observation to compute the forecasting error. We choose the optimal penalty in the ridge estimator such that the out-of-sample mean-squared forecast errors (MSFEs) achieve a minimum value. Thus, this setting is in line with the advocated models combining regularised estimation and data-driven selection of regularisation parameters in Abadie and Kasy (2019). In this experiment, we find that the optimal penalty $\widehat{\lambda}=0.8n$ for $h=1$ and $\widehat{\lambda}=0.7n$ for $h=2$.
@@
Next, for each configuration of $(\widehat{\lambda},h)=(0.8n,1)$ and $(0.7n,2)$, we execute Algorithm 1. We observe that the bias-correction procedure terminates within 3 to 5 iterations when setting $\eta=O(10^{-4})$ in equation (24). For simplicity, we opt for $k=10$ in the bias-correction, a choice deemed sufficiently large. Following the guidelines provided in Section 2.7, we compute pointwise prediction intervals for the data points in the testing subsample. The standard errors for each prediction are determined using the method described in Remark 5(iv). We present the $95\%$ pointwise prediction intervals for both 1-step ahead and 2-step ahead forecasts in Figure 7 and Figure 8, respectively. It is evident that the majority of the true observations fall within these prediction intervals. Additionally, we evaluate the coverage rates of the prediction intervals across the 138 points in the testing subsample, and find a consistent coverage rate of $92.03\%$ for both the 1-step and 2-step ahead forecasts, indicating that the $95\%$ prediction intervals perform effectively in this empirical example.
@@
Figure 9: Out-of-sample pointwise prediction intervals for the monthly CPI-All from July 1, 2008 to Dec. 12, 2019, constructed using (36)-(37) at $95\%$ level with $\widehat{\lambda}^{*}=\widehat{\lambda}=0.5n$, $k=10$, $n^{*}=31$, and $l=10$ for the $1$-step ahead forecast. The blue lines represent the confidence intervals, while the dark lines plot the true observations.